### v1

@@ Abstract
Inspired by the unconstrained PPE (UPPE) formulation [Liu, Liu, & Pego 2007 Comm. Pure Appl. Math., 60 pp. 1443], we previously proposed the GePUP formulation [Zhang 2016 J. Sci. Comput., 67 pp. 1134] for numerically solving the incompressible Navier-Stokes equations (INSE) on no-slip domains. In this paper, we propose GePUP-E and GePUP-ES, variants of GePUP that feature (a) electric boundary conditions with no explicit enforcement of the no-penetration condition, (b) equivalence to the no-slip INSE, (c) exponential decay of the divergence of an initially non-solenoidal velocity, and (d) monotonic decrease of the kinetic energy. Different from UPPE, the GePUP-E and GePUP-ES formulations are of strong forms and are designed for finite volume/difference methods under the framework of method of lines. Furthermore, we develop semi-discrete algorithms that preserve (c) and (d) and fully discrete algorithms that are fourth-order accurate for velocity both in time and in space. These algorithms employ algebraically stable time integrators in a black-box manner and only consist of solving a sequence of linear equations in each time step. Results of numerical tests confirm our analysis.
@@
Incompressible Navier-Stokes equations with no-slip conditions, Fourth-order accuracy, Projection methods, GePUP, Energy stability, Scalar auxiliary variable
pacs:
MSC Classification]Primary 76D05, 65M20
@@ 1 Introduction
The incompressible Navier-Stokes equations (INSE) with no-slip conditions read
@@
The INSE in (1) govern an enormous range of real-world phenomena such as blood flow, turbulence, atmosphere and ocean currents. As a fundamental and notoriously difficult problem, the well-posedness of the INSE (in $\mathbb{R}^{3}$ and $\mathbb{R}^{3}/\mathbb{Z}^{3}$ [1]) is selected by Smale [2] on his list of 18 math problems for the 21st century and by the Clay Institute on the list of 7 millennium problems [3]. Therefore, numerical computation appears to be the only viable way for now to obtain approximate solutions of the INSE. On the other hand, numerical computation may help the analytic understanding of INSE. As an example, Li [4] has recently shown that a bounded numerical solution with a small mesh size implies the existence of a smooth solution to the INSE.
1.1 Challenges for the design of numerical solvers
The aim of numerically solving differential equations is to efficiently, accurately, and faithfully reproduce the physical processes modeled by the equations. Specific to numerically solving the INSE are four major challenges that we confront in this work.
  1. [(A)]
  2. 1.
@@ 1.1 Challenges for the design of numerical solvers
(A) concerns the prominent feature of mass conservation: neither source nor sink exists anywhere inside the domain. It is well-known that a violation of (1b), even with small errors, might lead to qualitatively different flow patterns, especially for large Reynolds numbers. Another important physical constraint to be fulfilled in this work is the monotonic decrease of the total kinetic energy as defined in (51).
In (B), a crucial and indispensable type of numerical stability is the eigenvalue stability for the main evolutionary variable, which is typically the velocity. In addition, preserving the monotonicity of kinetic energy is equivalent to ensuring numerical stability on the velocity with respect to the 2-norm.
Challenge (C) concerns accuracy and efficiency. Near no-slip boundaries, flows at high Reynolds numbers tend to develop structures of multiple length scales and time scales. A numerical method should resolve all scales that are relevant to the important physics. Compared with fourth- and higher-order methods, first- and second-order methods have simpler algorithms and cheaper computations, but towards a given accuracy the computational resources may be rapidly exhausted. It is shown both theoretically and numerically in [5, Sec. 7] that fourth-order methods may have a large efficiency advantage over second-order methods.
Challenge (C) also concerns faithfully simulating flows where velocity derivatives such as vorticity crucially affect the physics. For first-order finite volume/difference methods, the computed velocity converges, but the vorticity does not, nor does the velocity gradient tensor. Consequently, the $O(1)$ error in $\nabla\mathbf{u}$ may lead to structures different from that of the original flow. In other words, it is not clear whether or not solutions of a first-order method have converged to the right physics. Similar suspicions apply to second-order methods for flows where second derivatives of the velocity are important.
Challenge (D) concerns versatility and user-friendliness of the numerical solver. To cater for the problem at hand, it is often desirable to change the time integrator from one to another. For example, flows with high viscosity are usually stiff while those with small viscosity are not; accordingly, an implicit time integrator should be used in the former case while an explicit one is usually suitable for the latter. If the internal details of a time integrator are coupled into the INSE solver in a boilerplate manner, it would be difficult and very inconvenient to change the time integrator; see also the discussion in the paragraph under (4). Hence a time integrator should be treated as a black box: for the ordinary differential equation (ODE) $\frac{\mathrm{d}U}{\mathrm{d}t}=f(U,t)$, we should only need to feed into the time integrator the initial condition $U^{n}$ and samples of $f$ at a number of time instances to get the solution $U^{n+1}$ from the black box.
@@
1.2 Some previous methods
In the original projection method independently proposed by Chorin [6] and Temam [7], the initial condition $\mathbf{u}^{n}\approx\mathbf{u}(t^{n})$ is first advanced to an auxiliary velocity $\mathbf{u}^{*}$ without worrying about the pressure gradient term and then $\mathbf{u}^{*}$ is projected to the divergence-free space to obtain $\mathbf{u}^{n+1}$,
@@ 1.2 Some previous methods
where $k$ is the time step size, $t^{n}$ is the starting time of the $n$th step, $\mathbf{g}^{n}\approx\mathbf{g}(t^{n})$, $\mathbf{C}(\mathbf{u}^{*},\mathbf{u}^{n})\approx[(\mathbf{u}\cdot\nabla)\mathbf{u}](t^{n})$, and ${\mathbf{L}}$ and ${\mathbf{P}}$ are discrete approximations of the Laplacian $\Delta$ and the Leray-Helmholtz projection ${\mathscr{P}}$, respectively; see Section 2.1.
@@ 1.2.1 Second-order methods with fractional time stepping
Chorin and Temam’s original projection method is first-order accurate and its improvement to the second order has been the aim of many subsequent works; see, e.g., [8, 9, 10, 11, 12, 13] and references therein. A common basis of many second-order methods is the temporal discretization of (1) with the trapezoidal rule,
@@
where $p^{n+\frac{1}{2}}\approx p(t^{n+\frac{1}{2}})$, $\mathbf{g}^{n+\frac{1}{2}}\approx\mathbf{g}(t^{n+\frac{1}{2}})$, and $[(\mathbf{u}\cdot\nabla)\mathbf{u}]^{n+\frac{1}{2}}\approx[(\mathbf{u}\cdot\nabla)\mathbf{u}](t^{n+\frac{1}{2}})$ are numerical approximations at the mid-time $t^{n+\frac{1}{2}}=\frac{1}{2}(t^{n}+t^{n+1})$ of the $n$th step.
[…]
@@
Since ${\mathbf{G}}^{T}=-{\mathbf{D}}$, the matrix $A$ has a saddle point structure and the above method is often called the saddle point approach. Despite its simplicity, this approach has two main disadvantages. First, the spatial discretization and time integration are coupled in a boilerplate manner and thus a change of either part would demand a complete rederivation of the matrix $A$. For fourth- or higher-order accuracy in time integration, it is often too complicated to have an explicit expression of the matrix $A$, as $A$ contains all internal details of the time integrator. Consequently, it is highly difficult for this approach to address challenges (C) and (D). Second, it is challenging [14] to efficiently solve the linear system (4) since all the velocity components and the pressure are coupled into a big unknown vector; in contrast, Chorin’s projection method only requires the solutions of linear systems with the unknowns as either the pressure or a velocity component.
Following (3), many second-order projection methods with fractional stepping can be united [13] as follows,
  • choose a boundary condition $\mathbf{B}(\mathbf{u}^{*})=\mathbf{0}$ to solve for $\mathbf{u}^{*}$,
    $\frac{\mathbf{u}^{*}-\mathbf{u}^{n}}{k}+\nabla q=-[(\mathbf{u}\cdot\nabla)\mathbf{u}]^{n+\frac{1}{2}}+\mathbf{g}^{n+\frac{1}{2}}+\frac{\nu}{2}\Delta(\mathbf{u}^{*}+\mathbf{u}^{n});$ (5)
  • enforce (3b) with a discrete projection so that
    $\mathbf{u}^{*}=\mathbf{u}^{n+1}+k\nabla\phi^{n+1},$ (6)
  • update the pressure by
    $p^{n+\frac{1}{2}}=q+\mathbf{U}(\phi^{n+1}),$ (7)
where the specific forms of $q$, $\mathbf{B}(\mathbf{u}^{*})$, and $\mathbf{U}(\phi^{n+1})$ vary from one method to another. For examples, the method of Bell, Colella & Glaz (BCG) [9] is specified by
$q=p^{n-\frac{1}{2}},\quad\mathbf{B}(\mathbf{u}^{*})=\mathbf{u}^{*}|_{\partial\Omega}=\mathbf{0},\quad\nabla p^{n+\frac{1}{2}}=\nabla p^{n-\frac{1}{2}}+\nabla\phi^{n+1},$ (8)
while that of Kim & Moin [8] by
$q$ $=0,\quad\mathbf{n}\cdot\mathbf{u}^{*}|_{\partial\Omega}=0,\quad{\bm{\tau}}\cdot\mathbf{u}^{*}|_{\partial\Omega}=k{\bm{\tau}}\cdot\nabla\phi^{n}|_{\partial\Omega},$ (9a)
$\nabla p^{n+\frac{1}{2}}$ $=\nabla\phi^{n+1}-\frac{\nu k}{2}\nabla\Delta\phi^{n+1}.$ (9b)
As observed by Brown, Cortez & Minion [13], the choices of $q$, $\mathbf{B}(\mathbf{u}^{*})$, and $\mathbf{U}(\phi^{n+1})$ are not independent. Indeed, substitute (6) into (5), subtract (3a), and we have
$\nabla p^{n+\frac{1}{2}}=\nabla q+\nabla\phi^{n+1}-\frac{\nu k}{2}\Delta\nabla\phi^{n+1}.$ (10)
Consequently, the pressure computed by (8) is only first-order accurate [13] because its pressure update formula is at an $O(k)$ discrepancy with (10). On the other hand, the pressure update formula (9b) complies with (10) due to the commutativity of Laplacian and the gradient, but (9b) and (6) imply $p^{n+\frac{1}{2}}=\phi^{n+1}-\frac{\nu}{2}\nabla\cdot\mathbf{u}^{*}$, which may deteriorate the pressure accuracy to the first order [13]. Nonetheless, both methods (8) and (9) are second-order accurate for the velocity.
@@
Fractional-stepping methods have been successful, but it is difficult to improve them to fourth- and higher-order accuracy. As discussed above, the choices of $q$, $\mathbf{B}(\mathbf{u}^{*})$, and $\mathbf{U}(\phi^{n+1})$ are coupled according to internal details of the time integrator. Consequently, switching from one time integrator to another calls for a new derivation. Furthermore, although appearing divorced, the velocity and the pressure are still implicitly coupled by the boundary condition of the auxiliary variable $\mathbf{u}^{*}$, with the coupling determined not by physics but still by internal details of the time integrator. Therefore, the approach of fractional-stepping methods is not suitable for tackling challenges (C) and (D).
1.2.2 The formulation of the pressure Poisson equation (PPE)
@@
(1a), (1c), (13), and the additional boundary condition $\nabla\cdot\mathbf{u}=0$ on $\partial\Omega$ are collectively called the PPE formulation of the INSE on no-slip domains [15]. In terms of computation, however, the PPE formulation has a decisive advantage over the original INSE. If (1a) is discretized in time with (1b) as a constraint, the resulting index-2 differential algebraic system may suffer from large order reductions [16]. In contrast, replacing the divergence-free constraint with the PPE avoids this difficulty. As such, the PPE formulation allows the time integrator to be treated as a black box and thus to be easily changed; indeed, the pressure is an implicit function of $\mathbf{u}$ and its interaction with $\mathbf{u}$ is completely decoupled from internal details of the time integrator. Also, there is no need to introduce nonphysical auxiliary variables. These advantages of the PPE formulation lead to quite a number of successful numerical methods [17, 15, 18, 19, 20, 21, 22].
Unfortunately, as observed by Liu, Liu & Pego [20], (1a) and (13a) yield
@@ 1.2.3 The formulation of unconstrained PPE (UPPE)
By applying the Leray-Helmholtz projection ${\mathscr{P}}$ to (11), Liu, Liu & Pego [20] obtained
@@
where the zero right-hand side (RHS) is added for stability reasons. The divergence of (15) and the second identity in (27) give
@@
which, by the maximum principle of the heat equation, dictates an exponential decay of a nonzero $\nabla\cdot\mathbf{u}$ in $\Omega$. A juxtaposition of (16) and (14) exposes a prominent advantage of (15) that any divergence residue is now well over control.
With the identity $\nabla(\nabla\cdot\mathbf{u})=\Delta({\mathcal{I}}-{\mathscr{P}})\mathbf{u}$ and the Laplace-Leray commutator $[\Delta,{\mathscr{P}}]:=\Delta{\mathscr{P}}-{\mathscr{P}}\Delta$ (see Section 2.2), Liu, Liu & Pego [20] rewrote (15) as
@@
Based on their UPPE formulation, the same authors developed a slip-corrected projection method [21] that is third-order accurate both in time and in space.
So far the UPPE formulation only concerns finite element methods based on the weak formulation. It is worthwhile to adapt UPPE to a strong formulation for the MOL discretization with finite volume or finite difference methods, as this may be a viable way to answer the challenges posed in Section 1.1. From (17) and the contents in Sections 2.2 and 2.3, a strong form of UPPE can be deduced [21] as
@@
The PPE (13) and the UPPE (19c,d) have slightly different forms and nonetheless a crucial distinction: it follows from the divergence of (19a) that (19c) leads to (16) whereas (13a) leads to (14).
Unfortunately, (19) is not yet suitable for the design of MOL-type finite volume and finite difference methods, due to two main reasons.
  1. [(RSN-1)]
  2. 1.
@@
1.2.4 The formulation of generic projection and unconstrained PPE (GePUP)

### v2

@@ Abstract
Inspired by the unconstrained PPE (UPPE) formulation [Liu, Liu, & Pego 2007 Comm. Pure Appl. Math., 60 pp. 1443], we previously proposed the GePUP formulation [Zhang 2016 J. Sci. Comput., 67 pp. 1134] for numerically solving incompressible Navier-Stokes equations (INSE) on no-slip domains. In this paper, we propose GePUP-E and GePUP-ES, variants of GePUP that feature (a) electric boundary conditions with no explicit enforcement of the no-penetration condition, (b) equivalence to the no-slip INSE, (c) exponential decay of the divergence of an initially non-solenoidal velocity, and (d) monotonic decrease of the kinetic energy. Different from UPPE, the GePUP-E and GePUP-ES formulations are of strong forms and are designed for finite volume/difference methods under the framework of method of lines. Furthermore, we develop semi-discrete algorithms that preserve (c) and (d) and fully discrete algorithms that are fourth-order accurate for velocity both in time and in space. These algorithms employ algebraically stable time integrators in a black-box manner and only consist of solving a sequence of linear equations in each time step. Results of numerical tests confirm our analysis.
@@
Incompressible Navier-Stokes equations with no-slip conditions Projection methods Energy stability Scalar auxiliary variable Fourth-order accuracy GePUP.
MSC:
Primary 76D05 65M20
journal:
@@ 1 Introduction
The incompressible Navier-Stokes equations (INSE) with no-slip conditions govern an enormous range of real-world phenomena such as blood flow, turbulence, atmosphere and ocean currents; they read
@@
In addition to efficiently, accurately, and faithfully reproducing the physical processes modeled by the equations, we confront four major challenges (FMC) of numerically solving the INSE,
  1. (FMC-1)
@@
(FMC-1) concerns the prominent feature of mass conservation: neither source nor sink exists anywhere inside the domain. It is well known that a violation of (1b), even with small errors, might lead to qualitatively different flow patterns, especially for large Reynolds numbers. Another important physical constraint to be fulfilled in this work is the monotonic decrease of the total kinetic energy as defined in (44).
In (FMC-2), a crucial and indispensable type of numerical stability is the eigenvalue stability for the main evolutionary variable, which is typically the velocity. In addition, preserving the monotonicity of kinetic energy is equivalent to ensuring numerical stability on the velocity with respect to the 2-norm.
Challenge (FMC-3) concerns accuracy and efficiency. Near no-slip boundaries, flows at high Reynolds numbers tend to develop structures of multiple length scales and time scales. A numerical method should resolve all scales that are relevant to the important physics. Compared with fourth- and higher-order methods, first- and second-order methods have simpler algorithms and cheaper computations, but towards a given accuracy the computational resources may be rapidly exhausted. It is shown both theoretically and numerically in (zhang16:_GePUP, , Sec. 7) that fourth-order methods may have a large efficiency advantage over second-order methods.
Challenge (FMC-3) also concerns faithfully simulating flows where velocity derivatives such as vorticity crucially affect the physics. For first-order finite volume/difference methods, the computed velocity converges, but the vorticity does not, nor does the velocity gradient tensor. Consequently, the $O(1)$ error in $\nabla\mathbf{u}$ may lead to structures different from that of the original flow. In other words, it is not clear whether or not solutions of a first-order method have converged to the right physics. Similar suspicions apply to second-order methods for flows where second derivatives of the velocity are important.
Challenge (FMC-4) concerns versatility and user-friendliness of the numerical solver. To cater for the problem at hand, it is often desirable to change the time integrator from one to another. For example, flows with high viscosity are usually stiff while those with small viscosity are not; accordingly, an implicit time integrator should be used in the former case while an explicit one is usually suitable for the latter. If the internal details of a time integrator are coupled into the INSE solver in a boilerplate manner, it would be difficult and very inconvenient to change the time integrator; see also the discussion in the paragraph under (4). Hence a time integrator should be treated as a black box: for the ordinary differential equation (ODE) $\frac{\mathrm{d}U}{\mathrm{d}t}=f(U,t)$, we should only need to feed into the time integrator the initial condition $U^{n}$ and samples of $f$ at a number of time instances to get the solution $U^{n+1}$ from the black box.
@@
1.1 Previous methods related to this work
In the original projection method independently proposed by Chorin Chorin_1968_numerical and Temam temam69:_sur_navier_stokes , the initial condition $\mathbf{u}^{n}\approx\mathbf{u}(t^{n})$ is first advanced to an auxiliary velocity $\mathbf{u}^{*}$ without worrying about the pressure gradient term and then $\mathbf{u}^{*}$ is projected to the divergence-free space to obtain $\mathbf{u}^{n+1}$,
@@ 1.1 Previous methods related to this work
where $k$ is the time step size, $t^{n}$ the starting time of the $n$th step, $\mathbf{g}^{n}\approx\mathbf{g}(t^{n})$, $\mathbf{C}(\mathbf{u}^{*},\mathbf{u}^{n})\approx[(\mathbf{u}\cdot\nabla)\mathbf{u}](t^{n})$, and $\mathbf{L}$} and $\mathbf{P}$} discrete approximations of the Laplacian $\Delta$ and the Leray-Helmholtz projection $\mathscr{P}$}, respectively; see Section 2.1.
@@ 1.1.1 Second-order methods with fractional time stepping
The original projection method is first-order accurate and its improvement to the second order has been the aim of many subsequent works; see, e.g., Kim.Moin_1985_application ; bell89:_secon_order_projec_method_incom ; orszag86:_bound_condit_incom_flows ; e.03:_gauge ; guermond06:_overview_of_projection_methods_for_incompressible_flows ; brown01:_accurate_projection_methods_for_ins and references therein. A common basis of many second-order methods is the temporal discretization of (1) with the trapezoidal rule,
@@
where $p^{n+\frac{1}{2}}\approx p(t^{n+\frac{1}{2}})$, $\mathbf{g}^{n+\frac{1}{2}}\approx\mathbf{g}(t^{n+\frac{1}{2}})$, and $(\mathbf{u}\cdot\nabla)\mathbf{u}]^{n+\frac{1}{2}}\approx[(\mathbf{u}\cdot\nabla)\mathbf{u}](t^{n+\frac{1}{2}})$ are numerical approximations at $t^{n+\frac{1}{2}}:=\frac{1}{2}(t^{n}+t^{n+1})$.
Replacing the gradient $\nabla$, the divergence $\nabla\cdot$, and the Laplacian $\Delta$ in (3) respectively with their second-order discrete counterparts $\mathbf{G}$}, $\mathbf{D}$}, and $\mathbf{L}$} yields
@@
Since ${\mathbf{G}}^{T}=-{\mathbf{D}}$, the matrix $A$ has a saddle point structure and the above method is often called the saddle point approach. Despite its simplicity, this approach has two main disadvantages. First, the spatial discretization and time integration are coupled in a boilerplate manner and thus a change of either part would demand a complete rederivation of the matrix $A$. For fourth- or higher-order accuracy in time integration, it is often too complicated to have an explicit expression of the matrix $A$, as $A$ contains all internal details of the time integrator. Consequently, it is highly difficult for this approach to address challenges (FMC-3,4). Second, it is challenging benzi05:_numer to efficiently solve the linear system (4) since all the velocity components and the pressure are coupled into a big unknown vector; in contrast, Chorin’s projection method only requires the solutions of linear systems with the unknowns as either the pressure or a velocity component.
In the fractional-stepping projection methods Kim.Moin_1985_application ; bell89:_secon_order_projec_method_incom ; brown01:_accurate_projection_methods_for_ins , one replaces $\mathbf{u}^{n+1}$ and $p^{n+\frac{1}{2}}$ in (3a) respectively with $\mathbf{u}^{*}$ and $q$, solves for the auxiliary velocity $\mathbf{u}^{*}$ with some boundary condition $\mathbf{B}(\mathbf{u}^{*})=\mathbf{0}$, obtains $\mathbf{u}^{n+1}$ by the projection $\mathbf{u}^{n+1}=\mathbf{u}^{*}-k\nabla\phi^{n+1}$, and updates the pressure with $p^{n+\frac{1}{2}}=q+\mathbf{U}(\phi^{n+1})$.
@@
Fractional-stepping projection methods have been very successful. However, the choices of $q$, $\mathbf{B}(\mathbf{u}^{*})$, and $\mathbf{U}(\phi^{n+1})$ are coupled according to internal details of the time integrator brown01:_accurate_projection_methods_for_ins . Consequently, switching from one time integrator to another calls for a new derivation. Furthermore, although appearing divorced, the velocity and the pressure are still implicitly coupled by the boundary condition of $\mathbf{u}^{*}$, with the coupling determined not by physics but still by internal details of the time integrator. Hence these methods are not suitable for tackling the challenges (FMC-3,4) either.
1.1.2 The formulation via the pressure Poisson equation (PPE)
@@
(1a), (1c), (7), and the additional boundary condition $\nabla\cdot\mathbf{u}=0$ on $\partial\Omega$ are collectively called the PPE formulation of the INSE on no-slip domains gresho87:_navier_stokes . In terms of computation, however, the PPE formulation has a decisive advantage over the original INSE. If (1a) is discretized in time with (1b) as a constraint, the resulting index-2 differential algebraic system may suffer from large order reductions sanderse12:_accur_runge_kutta_navier_stokes . In contrast, replacing the divergence-free constraint with the PPE avoids this difficulty. As such, the PPE formulation allows the time integrator to be treated as a black box and thus to be easily changed; indeed, the pressure is an implicit function of $\mathbf{u}$ and its interaction with $\mathbf{u}$ is completely decoupled from internal details of the time integrator. Also, there is no need to introduce nonphysical auxiliary variables. These advantages of the PPE formulation lead to quite a number of successful numerical methods kleiser80:_treat_incom_bound_condit_d ; gresho87:_navier_stokes ; henshaw94:_fourt_accur_method_incom_navier ; johnston04:_accur_navier_stokes ; liu07:_stabil_conver_effic_navier_stokes ; liu10_stable_accurate_pressure_unsteady_incompressible_viscous_flow ; shirokoff11:_navier_stokes .
Unfortunately, as observed by Liu, Liu & Pego liu07:_stabil_conver_effic_navier_stokes , (1a) and (7a) yield
@@ 1.1.3 The formulation of unconstrained PPE (UPPE)
The application of the Leray-Helmholtz projection $\mathscr{P}$} to (5) yields
@@
where the zero right-hand side (RHS) is added for stability reasons liu07:_stabil_conver_effic_navier_stokes . The divergence of (9) and the second identity in (20) give
@@
which, by the maximum principle of the heat equation, dictates an exponential decay of $\nabla\cdot\mathbf{u}$ in $\Omega$. A juxtaposition of (10) and (8) exposes a prominent advantage of (9) that any divergence residue is now well under control.
Via the identity $\nabla(\nabla\cdot\mathbf{u})=\Delta({\mathcal{I}}-{\mathscr{P}})\mathbf{u}$ and the Laplace-Leray commutator $\Delta,{\mathscr{P}}]:=\Delta{\mathscr{P}}-{\mathscr{P}}\Delta$, cf. Section 2.2, Liu, Liu & Pego liu07:_stabil_conver_effic_navier_stokes rewrote (9) as
@@
Based on the UPPE formulation, a slip-corrected projection method liu10_stable_accurate_pressure_unsteady_incompressible_viscous_flow is developed with third-order accuracy both in time and in space.
From (11) and the contents in Sections 2.2 and 2.3, a strong form of UPPE can be deduced liu10_stable_accurate_pressure_unsteady_incompressible_viscous_flow as
@@
The PPE (7) and the UPPE (13c,d) have slightly different forms and nonetheless a crucial distinction: it follows from the divergence of (13a) that (13c) leads to (10) whereas (7a) leads to (8).
Unfortunately, (13) is not yet suitable for the design of MOL-type finite volume and finite difference methods, due to two main reasons (TMR).
  1. (TMR-1)
@@
1.1.4 Generic projection and unconstrained PPE (GePUP)